OpenStudy (anonymous):

In store A the price of mayonnaise went down from $4.31 to $3.71. in store B it went from $ 8.09 to $7.23. In store C it went from $8.07 to 3.53 . what was the average decrease?

OpenStudy (anonymous):

add the orininal peize.. (1) find total decrease (2) now (1/2)*100 u will get it......

OpenStudy (anonymous):

do i add all the original prices together?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

does this mean 'average decrease as a price' or 'average decrease as a percent'?

OpenStudy (anonymous):

as a price

OpenStudy (anonymous):

okhay i added the original prices now what do i do?

OpenStudy (anonymous):

find the decrease in the price at each shop........

OpenStudy (anonymous):

now add all the decrease from each shop then to find average decrease divide the total decrease bay 3( since three terms )

OpenStudy (anonymous):

yea i got it wrong

OpenStudy (asadkarim7):

1st decrese = 4.31-3.71= 0.6 2nd decrese = 8.09-7.23=0.83 3rd decrese = 8.07-3.53=4.54 avg = sum of data / no of data 0.6 - 0.83-4.54 / 3 5.97/3= 1.99

OpenStudy (anonymous):

thanksssss

OpenStudy (asadkarim7):

is answer right?

OpenStudy (anonymous):

yes can you help me again?

OpenStudy (asadkarim7):

yeah sure